### v2

@@ 5.1 Local block likelihood
To evaluate the convolution in $l^{k,i-1}(\mathbf{x}_{i})$, we require all pixels in the set $\Phi_{i}$, which contains the pixels in block $i$ and all pixels within a frame of width $2r$ around block $i$. In the following, we denote this extended block by $\mathbf{x}_{i}^{+rr}\eqqcolon\mathbf{x}_{\Phi_{i}}\in\mathbb{R}^{q_{i}^{+rr}}$. Correspondingly, we let $\mathbf{x}_{i}^{+r}\in\mathbb{R}^{q_{i}^{+r}}$ denote the block $\mathbf{x}_{i}$ extended by all pixels within a frame of width $r$ around block $i$. Note that the extended blocks $\mathbf{x}_{i}^{+rr}$ and $\mathbf{x}_{i}^{+r}$ have different sizes $q_{i}^{+rr}$ and $q_{i}^{+r}$, respectively, based on their location in the image. For example, $\mathbf{x}_{i}^{+rr}$ has size $q_{i}^{+rr}\in\{(m+2r)^{2},(m+4r)(m+2r),(m+4r)^{2}\}$, depending on whether block $j$ is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrices $\mathbf{U}_{i}^{+r}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+r}}$ and $\mathbf{U}_{i}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}}$, which allow for the following mappings among $\mathbf{x}_{i}^{+rr}$, $\mathbf{x}_{i}^{+r}$, and $\mathbf{x}_{i}$:
${\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{x}_{i}^{+rr}=\mathbf{x}_{i}^{+r},\qquad{\mathbf{U}_{i}}^{\rm T}\mathbf{x}_{i}^{+rr}=\mathbf{x}_{i}.$
@@
(22) $l^{k,i-1}(\mathbf{x}_{i}^{+rr})=\frac{\lambda}{2}\|\mathbf{y}_{i}^{+r}-{\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{A}_{i}^{+rr}\mathbf{x}_{i}^{+rr}\|_{2}^{2},$
where $\mathbf{A}_{i}^{+rr}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+rr}}$ is the convolution matrix with the correct dimensions for the extended block $\mathbf{x}_{i}^{+rr}$. Moreover, we let
(23) $\mathbf{A}_{i}$ $={\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{A}_{i}^{+rr}\mathbf{U}_{i}$
(24) $\mathbf{y}^{k,i-1}$ $=\mathbf{y}_{i}^{+r}-{\mathbf{U}_{i}^{+r}}^{\rm T}\mathbf{A}_{i}^{+rr}\mathbf{P}_{i}^{+rr}\mathbf{x}_{i}^{+rr}$
where $\mathbf{P}_{i}^{+rr}=\mathbf{I}-\mathbf{U}_{i}{\mathbf{U}_{i}}^{\rm T}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+rr}}$ is an orthogonal projector, which projects $\mathbf{x}_{i}^{+rr}$ onto the fixed pixels of the current update, see the hatched area in Figure 2. Then, we can formulate Eq. 22 as a function in $\mathbf{x}_{i}$:
(25) $l^{k,i-1}(\mathbf{x}_{i})=\frac{\lambda}{2}\|\mathbf{y}^{k,i-1}-\mathbf{A}_{i}\mathbf{x}_{i}\|_{2}^{2},$
@@
(26) $\nabla l^{k,i-1}(\mathbf{x}_{i})=-\lambda\mathbf{A}_{i}^{\rm T}(\mathbf{y}^{k,i-1}-\mathbf{A}_{i}\mathbf{x}_{i}).$
@@ 5.2 Local block prior
The set of pixels $\Phi_{i}$ on which the conditional Eq. 20 depends is determined by the convolution. To compute the TV in $\varphi_{\varepsilon}^{k,i-1}$, we do not require all pixels in $\Phi_{i}$, but only the pixels of block $i$, extended by a frame of one pixel around it in the similar way as illustrated in Figure 2 with $12r=$ 1. In the following, we denote this extended block by $\mathbf{x}_{i}^{+1}\in\mathbb{R}^{q_{i}^{+1}}$. Note that the extended block $\mathbf{x}_{i}^{+1}$ can have different sizes $q_{i}^{+1}$, depending on whether it is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrix $\mathbf{W}_{i}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}}$, which allows for the following mapping between $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{+1}$:
${\mathbf{W}_{i}}^{\rm T}\mathbf{x}_{i}^{+1}=\mathbf{x}_{i}.$
@@
$\mathbf{D}_{i}^{l}={\mathbf{D}_{i}^{l}}^{+1}\mathbf{W}_{i},\qquad\mathbf{b}_{l}^{k,i-1}={\mathbf{D}_{i}^{l}}^{+1}\mathbf{P}_{i}^{+1}\mathbf{x}_{i}^{+1},$
where ${\mathbf{D}_{i}^{l}}^{+1}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}^{+1}}$ computes the differences in vertical ($l=v$) or horizontal ($l=h$) direction of the extended block $\mathbf{x}_{i}^{+1}$. Moreover, the orthogonal projector $\mathbf{P}_{i}^{+1}=\mathbf{I}-\mathbf{W}_{i}{\mathbf{W}_{i}}^{\rm T}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}^{+1}}$ projects $\mathbf{x}_{i}^{+1}$ onto the pixels, which are fixed during the update, see the hatched area in Figure 2. Then, we can evaluate the potential of the local block prior by
(27) $\varphi_{\varepsilon}^{k,i-1}(\mathbf{x}_{i})=\delta\sum_{\alpha}\sqrt{[\mathbf{D}_{i}^{v}\mathbf{x}_{i}+\mathbf{b}_{v}^{k,i-1}]_{\alpha}^{2}+[\mathbf{D}_{i}^{h}\mathbf{x}_{i}+\mathbf{b}_{h}^{k,i-1}]_{\alpha}^{2}+\varepsilon},$
@@
(28) $\nabla\varphi_{\varepsilon}^{k,i-1}(\mathbf{x}_{i})={\mathbf{D}_{i}^{v}}^{\rm T}\Lambda(\mathbf{x}_{i})^{-1}(\mathbf{D}_{i}^{v}\mathbf{x}_{i}+\mathbf{b}_{v}^{k,i-1})+{\mathbf{D}_{i}^{h}}^{\rm T}\Lambda(\mathbf{x}_{i})^{-1}(\mathbf{D}_{i}^{h}\mathbf{x}_{i}+\mathbf{b}_{h}^{k,i-1}),$
with
$\Lambda(\mathbf{x}_{i})=\mathrm{diag}\left(\sqrt{[\mathbf{D}_{i}^{v}\mathbf{x}_{i}+\mathbf{b}_{v}^{k,i-1}]_{\alpha}^{2}+[\mathbf{D}_{i}^{h}\mathbf{x}_{i}+\mathbf{b}_{h}^{k,i-1}]_{\alpha}^{2}+\varepsilon}\right).$
@@ 5.3 Local & parallel MLwG algorithm
Due to the local conditional dependencies in the posterior, we may be able to update several blocks in parallel during the loop in line 3 in Algorithm 1. However, when updating block $i$, we have to fix all pixels within a frame of width $2r$ around this block, and therefore, we can not update any of the neighboring blocks at the same time. For an example, see Figure 1, where the hatched area corresponds to the fixed blocks when updating block $\mathbf{x}_{6}$.
We define a parallel updating scheme via the index sets $\mathcal{U}_{l}\subset[b],\,l=1,2,\dots$ with $\mathcal{U}_{l}\cap\mathcal{U}_{p}=\varnothing~{}\forall l\neq p$, such that the blocks $\{i\,|\,i\in\mathcal{U}_{l}\}$ can be updated in parallel. The choices $\mathcal{U}_{l}$ are not unique, but it holds $l\geq 4$ for 2D images with square block partition. In this paper, we use the minimal number of required updating sets $\mathcal{U}_{l}$, i.e., $l\in\{1,2,3,4\}$. With this choice, the number of blocks which can be updated in parallel is
@@
In Algorithm 2, we state a local and parallel version of the MLwG Algorithm 1. That is, we consider parallel updates in line 4, and we evaluate the full conditionals in the block updates locally, i.e., we use the expressions for the local block likelihood Eq. 25 and the local block prior Eq. 27 in line 7.
@@
0:  Number of samples $N$, step size $\tau$, initial state $\mathbf{x}^{0}$, partition into blocks $i=1,\dots,b$, sets of indices for parallel updating $\mathcal{U}_{l}$.
@@
5:           Set $\mathbf{x}^{k,i}\leftarrow\mathbf{x}^{k,i-1}$.
6:           Get $\mathbf{x}_{i}^{+rr}$ and $\mathbf{x}_{i}^{+1}$ from the current state $\mathbf{x}^{k,i}$ (see Fig. 2 for example).
7:           Propose a block candidate $\mathbf{z}_{i}^{k,i}$ via Eq. 12 for the local block density Eq. 21, with potentials given by Eqs. 25, LABEL: and 27. The gradients of the potentials are given by Eqs. 26, LABEL: and 28.
@@ 6 Numerical examples
Both images in the following experiments are in grayscale and are normalized such that the pixel values are between 0 and 1. Note however, that we do not enforce the box constraint $[0,1]$ on the pixel values in our experiments. In each sampling experiment, we compute 5 independent sample chains with $2000$ samples each and apply thinning by saving only every $200$-th sample to reduce correlation. We check our sample chains for convergence by means of the potential scale reduction factor (PSRF) [10]. In brief, PSRF compares the within-variance with the in-between variance of the chains. Empirically, one considers sample chains to be converged if $\mathrm{PSRF}<1.1$. We use the Python package arviz [14] to compute the PSRF. With the same package, we compute the normalized effective sample size (nESS) and credibility intervals (CI), see, e.g., [21] for definitions.
@@ 6.1.1 Problem description
We consider a blurred and noisy image, “cameraman”, with the size $512\times 512$, and show the “true” image on the left in Figure 3. The different image sections defined by the black frames and block partitions in the same image are required for the experiments in Sections 6.1.3, LABEL: and 6.1.4. The data is synthetically obtained via the observation model Eq. 3, where $\mathbf{A}\in\mathbb{R}^{d\times d}$ corresponds to the discretization of a Gaussian blurring kernel with radius $8$ and standard deviation $8$. The noise is a realization of $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},0.01^{2}\mathbf{I})$, and the degraded “cameraman” is shown on the right in Figure 3.
@@ 6.1.2 Influence of the smoothing parameter $\bm{\varepsilon}$
We compute MAP estimates for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$ with the majorization-minimization algorithm proposed in [8] and show the results in the left column of Figure 4. We can see that the restoration from $\varepsilon=10^{-3}$ is less cartoon-like and has smoother edges than the other two restorations. The difference between the results from $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ is hardly visible.
@@
We now run our local & parallel MLwG (Algorithm 2) for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$ with step size adaptation during burn-in [19] targeting an acceptance rate of $0.547$ in each block. We chose this target acceptance rate due to the results in [29]. We show the sample means and widths of the 90% sample CIs in Figure 4. Similar to the MAP estimates, the sample means of $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ are more favorable than the result of $\varepsilon=10^{-3}$ because it contains visible artifacts.
Moreover, the 90% sample CI difference is in general wider for $\varepsilon=10^{-3}$ than for $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$. However, on the edges the width of the 90% sample CIs are rather similar. We also list some results about the sample chains in Table 1. Here we note that $\varepsilon=10^{-3}$ allows for a significantly larger step size in comparison to $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$. This results in less correlated samples, which is reflected in a larger nESS.
We conclude that relatively small values of $\varepsilon$ make the posterior density smoother, allowing for larger step sizes, and thus making the sampling more efficient in terms of nESS. However, at the same time, the results can be visually significantly different compared to choices of small $\varepsilon$, which yield sharper edges in the MAP estimate and the mean. Based on these observations, and since the results for $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ are very close, we fix $\varepsilon=10^{-5}$ in the remaining experiments.
Table 1: Sampling results of MLwG for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$. The shown nESS is the mean over all samples. The shown step size $\tau$ and acceptance rate $\alpha$ are the mean over all blocks. The maximum and median PSRF are with respect to all pixels.
$\bm{\varepsilon}$ nESS [%] $\bm{\tau}$ [10-6] $\bm{\alpha}$ [%] max PSRF median PSRF
$10^{-3}$ 53.5 25.8 54.7 1.01 1.00
$10^{-5}$ 27.0 7.5 54.4 1.03 1.00
$10^{-7}$ 21.2 5.6 54.3 1.04 1.00

### v3

@@ 5.1 Local block likelihood
To evaluate the convolution in $l^{k,i-1}({\bm{x}}_{i})$, we require all pixels in the set $\Phi_{i}$, which contains the pixels in block $i$ and all pixels within a frame of radius $2r$ around block $i$. In the following, we denote this extended block by $\bm{x}_{i}^{+rr}\eqqcolon{\bm{x}}_{\Phi_{i}}\in\mathbb{R}^{q_{i}^{+rr}}$. Correspondingly, we let $\bm{x}_{i}^{+r}\in\mathbb{R}^{q_{i}^{+r}}$ denote the block $\bm{x}_{i}$ extended by all pixels within a frame of radius $r$ around block $i$. Note that the extended blocks $\bm{x}_{i}^{+rr}$ and $\bm{x}_{i}^{+r}$ have different sizes $q_{i}^{+rr}$ and $q_{i}^{+r}$, respectively, based on their location in the image. For example, $\bm{x}_{i}^{+rr}$ has size $q_{i}^{+rr}\in\{(m+2r)^{2},(m+4r)(m+2r),(m+4r)^{2}\}$, depending on whether block $i$ is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrices $\bm{U}_{i}^{+r}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+r}}$ and $\bm{U}_{i}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}}$, which allow for the following mappings among $\bm{x}_{i}^{+rr}$, $\bm{x}_{i}^{+r}$, and $\bm{x}_{i}$:
${\bm{U}_{i}^{+r}}^{\rm T}\bm{x}_{i}^{+rr}=\bm{x}_{i}^{+r},\qquad{\bm{U}_{i}}^{\rm T}\bm{x}_{i}^{+rr}=\bm{x}_{i}.$
@@
(22) $l^{k,i-1}(\bm{x}_{i}^{+rr})=\frac{\lambda}{2}\|\bm{y}_{i}^{+r}-{\bm{U}_{i}^{+r}}^{\rm T}\bm{A}_{i}^{+rr}\bm{x}_{i}^{+rr}\|_{2}^{2},$
where $\bm{A}_{i}^{+rr}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+rr}}$ is the convolution matrix with the correct dimensions for the extended block $\bm{x}_{i}^{+rr}$. Moreover, we let
(23) $\bm{A}_{i}$ $={\bm{U}_{i}^{+r}}^{\rm T}\bm{A}_{i}^{+rr}\bm{U}_{i}$
(24) ${\bm{y}}^{k,i-1}$ $=\bm{y}_{i}^{+r}-{\bm{U}_{i}^{+r}}^{\rm T}\bm{A}_{i}^{+rr}\bm{P}_{i}^{+rr}\bm{x}_{i}^{+rr}$
where $\bm{P}_{i}^{+rr}=\bm{I}-\bm{U}_{i}{\bm{U}_{i}}^{\rm T}\in\mathbb{R}^{q_{i}^{+rr}\times q_{i}^{+rr}}$ is an orthogonal projector, which projects $\bm{x}_{i}^{+rr}$ onto the fixed pixels of the current update, see the hatched area in Figure 2. Then, we can formulate Eq. 22 as a function in $\bm{x}_{i}$:
(25) $l^{k,i-1}(\bm{x}_{i})=\frac{\lambda}{2}\|{\bm{y}}^{k,i-1}-\bm{A}_{i}\bm{x}_{i}\|_{2}^{2},$
@@
(26) $\nabla l^{k,i-1}({\bm{x}}_{i})=-\lambda\bm{A}_{i}^{\rm T}({\bm{y}}^{k,i-1}-\bm{A}_{i}\bm{x}_{i}).$
@@ 5.2 Local block prior
The set of pixels $\Phi_{i}$ on which the conditional Eq. 20 depends is determined by the convolution. To compute the TV in $\varphi_{\varepsilon}^{k,i-1}$, we do not require all pixels in $\Phi_{i}$, but only the pixels of block $i$, extended by a frame of one pixel around it. This is similar to the computation of the local block likelihood with $12r=$ 1, see Figure 2 again. In the following, we denote this extended block by $\bm{x}_{i}^{+1}\in\mathbb{R}^{q_{i}^{+1}}$. Note that the extended block $\bm{x}_{i}^{+1}$ can have different sizes $q_{i}^{+1}\in\{(m+1)^{2},(m+1)(m+2),(m+2)^{2}\}$, depending on whether it is located in a corner, at an edge or in the interior of the image. Moreover, we define the component selection matrix $\bm{W}_{i}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}}$, which allows for the following mapping between $\bm{x}_{i}$ and $\bm{x}_{i}^{+1}$:
${\bm{W}_{i}}^{\rm T}\bm{x}_{i}^{+1}=\bm{x}_{i}.$
@@
$\bm{D}_{i}^{(l)}={\bm{D}_{i}^{(l)}}^{+1}\bm{W}_{i},\qquad\bm{b}_{(l)}^{k,i-1}={\bm{D}_{i}^{(l)}}^{+1}\bm{P}_{i}^{+1}\bm{x}_{i}^{+1},$
where ${\bm{D}_{i}^{(l)}}^{+1}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}^{+1}}$ computes the differences in vertical ($l=v$) or horizontal ($l=h$) direction of the extended block $\bm{x}_{i}^{+1}$. Moreover, the orthogonal projector $\bm{P}_{i}^{+1}=\bm{I}-\bm{W}_{i}{\bm{W}_{i}}^{\rm T}\in\mathbb{R}^{q_{i}^{+1}\times q_{i}^{+1}}$ projects $\bm{x}_{i}^{+1}$ onto the pixels, which are fixed during the update. This is similar to the computation of the local block likelihood, see the hatched area in Figure 2 again. Then, we can evaluate the potential of the local block prior by
(27) $\varphi_{\varepsilon}^{k,i-1}(\bm{x}_{i})=\delta\sum_{\alpha}\sqrt{[\bm{D}_{i}^{(v)}\bm{x}_{i}+\bm{b}_{(v)}^{k,i-1}]_{\alpha}^{2}+[\bm{D}_{i}^{(h)}\bm{x}_{i}+\bm{b}_{(h)}^{k,i-1}]_{\alpha}^{2}+\varepsilon},$
@@
(28) $\begin{split}\nabla\varphi_{\varepsilon}^{k,i-1}({\bm{x}}_{i})=~{}&{\bm{D}_{i}^{(v)}}^{\rm T}\Lambda(\bm{x}_{i})^{-1}(\bm{D}_{i}^{(v)}\bm{x}_{i}+\bm{b}_{(v)}^{k,i-1})\\ &+{\bm{D}_{i}^{(h)}}^{\rm T}\Lambda(\bm{x}_{i})^{-1}(\bm{D}_{i}^{(h)}\bm{x}_{i}+\bm{b}_{(h)}^{k,i-1}),\\ \text{where}\quad\Lambda(\bm{x}_{i})={\rm diag}&\left(\sqrt{[\bm{D}_{i}^{(v)}\bm{x}_{i}+\bm{b}_{(v)}^{k,i-1}]_{\alpha}^{2}+[\bm{D}_{i}^{(h)}\bm{x}_{i}+\bm{b}_{(h)}^{k,i-1}]_{\alpha}^{2}+\varepsilon}\right).\end{split}$
Remark 4.
It is straightforward to extend the above computations to other types of difference operators (e.g., rank-deficient difference matrices, higher-order finite differences, etc.) as long as they exhibit the local structure, i.e., they act spatially local on the pixels.
@@ 5.3 Local & parallel MLwG algorithm
Due to the local conditional dependencies in the posterior, we may be able to update several blocks in parallel during the loop in line 3 in Algorithm 1. However, when updating block $i$, we have to fix all pixels within a frame of radius $2r$ around this block, and therefore, we can not update any of the neighboring blocks at the same time. For an example, see Figure 1, where the hatched area corresponds to the fixed blocks when updating block ${\bm{x}}_{6}$.
We define a parallel updating scheme via the index sets $\mathcal{U}_{l}\subset[b],\,l=1,2,\dots$ with $\mathcal{U}_{l}\cap\mathcal{U}_{p}=\varnothing,~{}\forall l\neq p$, such that the blocks $\{i\,|\,i\in\mathcal{U}_{l}\}$ can be updated in parallel. The choices $\mathcal{U}_{l}$ are not unique, but it holds $l\geq 4$ for 2D images with square block partition. In this paper, we use the minimal number of required updating sets $\mathcal{U}_{l}$, i.e., $l\in\{1,2,3,4\}$. With this choice, the number of blocks which can be updated in parallel is
@@
In Algorithm 2, we state a local & parallel version of the MLwG Algorithm 1. That is, we consider parallel updates in line 4, and we evaluate the full conditionals in the block updates locally, i.e., we use the expressions for the local block likelihood Eq. 25 and the local block prior Eq. 27 in line 7.
@@
0:  Number of samples $N$, step size $\tau$, initial state ${\bm{x}}^{0}$, partition into blocks $i=1,\dots,b$, sets of indices for parallel updating $\mathcal{U}_{l}$.
@@
5:           Set ${\bm{x}}^{k,i}\leftarrow{\bm{x}}^{k,i-1}$.
6:           Get $\bm{x}_{i}^{+rr}$ and $\bm{x}_{i}^{+1}$ from the current state ${\bm{x}}^{k,i}$ (see Fig. 2 for example).
7:           Propose a block candidate ${\bm{z}}_{i}^{k,i}$ via Eq. 12 for the local block density Eq. 21, with potentials given by Eqs. 25, LABEL: and 27. The gradients of the potentials are given by Eqs. 26, LABEL: and 28.
@@ 6 Numerical examples
Both images in the following experiments are in grayscale and are normalized such that the pixel values are between 0 and 1. Note however, that we do not enforce the box constraint $[0,1]$ on the pixel values in our experiments. In each sampling experiment, we compute 5 independent sample chains with $2000$ samples each and apply thinning by saving only every $200$-th sample to reduce correlation. We check our sample chains for convergence by means of the potential scale reduction factor (PSRF) [11]. In brief, PSRF compares the within-variance with the in-between variance of the chains. Empirically, one considers sample chains to be converged if $\mathrm{PSRF}<1.1$. We use the Python package arviz [15] to compute the PSRF. With the same package, we compute the normalized effective sample size (nESS) and credibility intervals (CI), see, e.g., [22] for definitions. Our code is publicly available in [9].
@@ 6.1.1 Problem description
We consider a blurred and noisy image, “cameraman”, with the size $512\times 512$ and show the “true” image on the left in Figure 3. The different image sections defined by the black frames and block partitions in the same image are required for the experiments in Sections 6.1.3, LABEL: and 6.1.4. The data is synthetically obtained via the observation model Eq. 3, where $\bm{A}\in\mathbb{R}^{d\times d}$ corresponds to the discretization of a Gaussian blurring kernel with radius $8$ and standard deviation $8$. The noise is a realization of $\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\bm{I})$, and the degraded “cameraman” is shown on the right in Figure 3.
@@ 6.1.2 Influence of the smoothing parameter $\bm{\varepsilon}$
We compute MAP estimates for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$ with the majorization-minimization algorithm proposed in [8] and show the results in the left column of Figure 4.
We can see that the restoration from $\varepsilon=10^{-3}$ has smoother edges than the other two restorations which exhibit the typical cartoon-like structure of TV-regularized images. Further, the difference between the results from $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ is hardly visible.
@@
We now run our local & parallel MLwG (Algorithm 2) for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$ with a diminishing step size adaptation during burn-in [20] targeting an acceptance rate of $0.547$ in each block.  [^1] We note that the diminishing step size adaptation proposed in [20] is performed during the whole sampling process. In our experiments, we stop the adaptation after a specified number of burn-in samples. We chose this target acceptance rate due to the results in [31]. We show the sample means and widths of the 90% sample CIs in Figure 4. Here, the sample means of $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ are visually more favorable than their corresponding MAP estimates. In contrast, the result of $\varepsilon=10^{-3}$ contains visible artifacts.
Moreover, the 90% sample CI difference is in general wider for $\varepsilon=10^{-3}$ than for $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$. However, on the edges, the width of the 90% sample CIs are rather similar. We also list some results about the sample chains in Table 1. Here, we note that $\varepsilon=10^{-3}$ allows for a significantly larger mean step size in comparison to $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$. This results in less correlated samples, which is reflected in a larger nESS.
We conclude that relatively large values of $\varepsilon$ make the posterior density smoother, allowing for larger step sizes and thus making the sampling more efficient in terms of nESS. However, at the same time, the results can be visually significantly different compared to choices of small $\varepsilon$, which yield sharper edges in the MAP estimate and the mean. Based on these observations, and since the results for $\varepsilon=10^{-5}$ and $\varepsilon=10^{-7}$ are very close, we fix $\varepsilon=10^{-5}$ in the remaining experiments.
Table 1: Sampling results of MLwG for $\varepsilon\in\{10^{-3},10^{-5},10^{-7}\}$. The shown min nESS is the pixel-wise minimum of the mean nESS which is taken over the $5$ chains. The shown step size $\tau$ and acceptance rate $\alpha$ are the means taken over all blocks and the $5$ chains. The maximum and median PSRF are with respect to the pixels.
$\bm{\varepsilon}$ min nESS [%] $\bm{\tau}$ [10-6] $\bm{\alpha}$ [%] max PSRF median PSRF
$10^{-3}$ 13.3 25.8 54.7 1.01 1.00
$10^{-5}$ 3.2 7.5 54.4 1.03 1.00
$10^{-7}$ 2.1 5.6 54.3 1.04 1.00